OpenStudy (anonymous):

Solve the equation 4/5y + 2 = 6 for y. Please show your work and solution in the space provided.

OpenStudy (mertsj):

\[\frac{4}{5y+2}=\frac{6}{1}\]

OpenStudy (mertsj):

Is that the correct problem?

OpenStudy (anonymous):

i have no idea what you just said but its 4/5 y + 2 = 6

OpenStudy (mertsj):

4/(5y+2)=6 pr 4/5y+2=6?

OpenStudy (mertsj):